A bicyclist rode into the country for 5hours. In returning, her speed was 5mi/h faster and the trip took 4 hours. What was her speed each way?
Answer by checkley75(3666)

5R=4(R+4)
5R=4R+16
5R-4R=16
R=16 IS THE RATE OF THE 5 HOUR TRIP WHICH WAS 5*16=80 MILES LONG
THUS THE 4 HOUR TRIP WAS 4(16+4)=80 MILES ALSO